# BootsTAP: Bootstrapped Training for Tracking-Any-Point

### v1

@@ 4 Experiments
We train our model on over 15 million 24-frame clips from from publicly-available online videos, in conjunction with standard training on Kubric. The resulting model is essentially a drop-in replacement for TAPIR (albeit with slightly larger computational requirements due to the extra layers). We evaluate on the TAP-Vid benchmark using the standard protocol.
@@ 4.1 Training datasets
We collected a video dataset from publicly accessible videos selected from categories that typically contain high-quality and realistic motion (such as lifestyle and one-shot videos). Conversely, we omitted videos from categories with low visual complexity or unrealistic motions, such as tutorial videos, lyrics videos, and animations. To maintain consistency, we exclusively obtained videos with shot at 60fps. Additionally, we applied a quality metric by only considering videos with over 200 views. We removed the first and last 2 seconds of each video, as these often contain intros and outros with text or other overlays. From each video, we randomly sampled five clips, excluding those with overlay/watermarked frames, which were identified by checking the horizontal and vertical gradients and computing the pixel-wise median (similar to [6]). Furthermore, we expect the teacher signal will be more reliable on continuous shots due to temporal continuity; therefore, clips with shot boundary changes are detected and removed based on [3, 62, 51, 35] with additional accuracy improvements based on full-frame geometric alignment. In total, we generated 15 million clips for training.
@@ 4.2 Evaluation datasets
We rely on the TAP-Vid [9] and RoboTAP [52] benchmarks for quantitative evaluation; in all cases, we evaluate zero-shot on the entire benchmark, resizing to $256\times 256$ before evaluating according to the standard procedure [9]. All evaluation datasets consists of real world videos.
TAP-Vid-Kinetics contains videos collected from the Kinetics-700-2020 validation set [5] with original focus on video action recognition. This benchmark contains 1K YouTube videos of diverse action categories, approximately 10 seconds long, including many challenging elements such as shot boundaries, multiple moving objects, dynamic camera motion, cluttered background and dark lighting conditions. Each video contains $\sim$26 tracked points on average, obtained from careful human annotation.
RoboCAT-NIST is a subset of the data collected for RoboCat [4]. Inspired by the NIST benchmark for robotic manipulation [26], it includes gears of varying sizes (small, medium, large) and a 3-peg base, introduced for a systematic study of insertion affordance. All videos are collected by human teleoperation. It includes 100,000 real world videos with different robot arms operating and inserting gears, which are a particularly challenging case due to the rotational symmetry and lack of texture. In this work, we processed videos to 64 frames long with 222x296 resolution. This dataset is mainly for demonstration purpose, there are no human groundtruth point tracks.
4.3 Evaluation metrics
For TAP-Vid datasets, evaluation is split into strided mode and first mode. Strided mode samples query points every 5 frames on the groundtruth tracks when they are visible. Query points can be any time in the video hence it tests the model prediction power both forward and backward in time. First mode samples query points only when they are first time visible and the evaluation only measures tracking accuracy in future frames.
@@ 4.4 Results
Table 2 shows performance under “first” metrics. Here, we see that bootstrapping outperforms prior works by a wide margin on Kinetics; this is likely because TAPIR’s global search is more robust to large occlusions and cuts, which are more prominent in Kinetics. This search might harm performance in datasets like DAVIS with a stronger temporal continuity bias. Perhaps most surprising is the very strong improvement in RoboTAP–almost 10% absolute performance–despite RoboTAP looking very different from typical online videos. We see similar results for RGB-Stacking in 1. These two datasets have large textureless regions; such regions are challenging to track without priors for object segmentation, which are difficult to obtain from synthetic datasets.
Figure 3: Comparison between TAPIR and BootsTAP on TAP-Vid-DAVIS. Top: we show query points, and below, we show predictions for the two methods for two later frames. Filled circles show un-occluded predictions relative to ground truth (GT) (ends of the associated segments). x’s indicate predictions where the GT is occluded, while empty circles are points visible in GT but predicted occluded. Below we show an alternative view of the same ground truth and predictions, where the entire trajectory is plotted over the entire frame to show the full motion. On challenging, cluttered scenes like these, BootsTAP significantly improves occlusion prediction accuracy (note the missing tracks on the motorbike and the people behind the dancer). It can also improve localization accuracy: note the large position errors from TAPIR on the motorbike ignoring the occlusion error.
Figure 3 shows two representative examples where BootsTAP improves TAPIR on TAP-Vid-DAVIS. We notice that the main improvements come from occlusion prediction and precise localization. The scooter example clearly shows BootsTAP outperforms on occlusion estimation, but also note that for some of the points TAPIR incorrectly marks as occluded, the localization estimate is also quite bad (the green and yellow points, predicted to be on background). In the breakdance example, we see that TAPIR marks several background points as occluded despite very little motion, likely due to the large number of people (note Kubric contains no humans). The improved occlusion accuracy here suggests that the model has learned about distinguishing different people from its experience with real videos.
Figure 4: Comparison between TAPIR and BootsTAP on RoboTAP dataset. We show the TAPIR prediction, BootsTAP prediction and ground truth tracks in Rainbow tail style in 3 columns respectively. The groundtruth tracks are obtained through human annotation [52]. Comparing to TAPIR, BootsTAP works more accurately on occlusion prediction in the real robotics scenes with camera motion. Particularly points that were failed in TAPIR can now be captured in BootsTAP.
Figure 4 shows more examples for where BootsTAP improves TAPIR on RoboTAP dataset. We show the groundtruth tracks, TAPIR prediction and BootsTAP prediction, where each trajectory is plotted over the last video frame with a different color. Even though RoboTAP contains different views, we select mainly the wrist camera views as examples where the improvements from BootsTAP are particularly clear. TAPIR loses many points due to the large motion and lack of texture, but BootsTAP recovers them.
@@
Figure 5: Comparison between TAPIR and BootsTAP on the real RoboCAT-NIST dataset. Due to no groundtruth, we show the TAPIR prediction and BootsTAP prediction in Rainbow tail style side-by-side. On NIST, BootsTAP works more consistently on location prediction. Particularly points that were originally predicted as occluded now can be visible.
@@
Figure 5 further illustrates improvements on the RoboCAT-NIST dataset, using a model where 100K RoboCAT-NIST clips were added to the online data. Due to the lack of groundtruth, we instead sample a grid of points on the gears (i.e. red pixels). We display a few examples comparing the predicted tracks between the two models. As these are rigid objects, we expect the points to move consistently within each gear; deviations from this are errors. Due to the lack of texture on the gears and the nontrivial sim2real transfer gap, the original TAPIR trained on Kubric works poorly here, with many jittery tracks and severe tracking failures. This is particularly bad for points that are close to occlusion or out of image boundary. The bootstrapped model fixes many of these failures: the tracks are much smoother and occlusion predictions become much more accurate.
@@ 5 Conclusion
Acknowledgements
. We thank Ignacio Rocco, Yusuf Aytar, David Fleet, Daniel Zoran, Jon Scholz, and Nando de Freitas for helpful conversations and advice. We thank Lucas Smaira, Ed Hirst, Alexandre Moufarek, Jack Parker-Holder and Junlin Zhang for help with dataset construction.
@@ Appendix 0.A Summary of the approach
We summarize notations and computation of our self-supervised loss in Algorithm 1.
Algorithm 1 BootsTAP self-supervised loss. Notations:
@@
       sample a point $Q_{2}=(q_{2},t_{2})$ from $\mathcal{U}(\mathcal{V})$.
    Else:
       Leave the point unchanged: $Q_{2}\leftarrow Q1$.
@@ Appendix 0.B Implementation details
0.B.1 Distribution over affine transformation
To enforce equivariance to spatial transformations, for the frames and query point coordinates, we apply a frame-wise affine transformation $\Phi=\{\Phi_{t}\}_{t}$, whose scaling parameters $(h_{t},w_{t})$ and translation parameters $(c_{t}^{x},c_{t}^{y})$ vary linearly over time, randomly sampled from a distribution $\mathcal{D}_{\Phi}$ that we describe next.
We first sample a pair of initial and final spatial dimensions $(H_{0},W_{0})$ and $(H_{T-1},W_{T-1})$ from a distribution $\mathcal{D}_{H,W}$ defined as follows. For each pair, we first sample an area $A$ uniformly over $[0.6,1.0]$. Next we sample values $a^{1},a^{2}\sim\mathcal{U}([A,1])$ and derive random height value by averaging them $h=\frac{a^{1}+a^{2}}{2}$ and width value $w=\frac{A}{h}$; and finally, we multiply these values by the input’s original shape $(H,W)$. This gives us a pair of spatial dimensions biased towards aspect ratios close to 1, and covering an area between $60\%$ and $100\%$ of the original input.
We also sample a pair of initial spatial coordinates $(C_{0}^{x},C_{0}^{y})$ for the top-left corner of the transformed view, within the frame, uniformly over $[0,P_{x})\times[0,P_{y})$, where $P_{x}=W-W_{0}$ and $P_{y}=H-H_{0}$ respectively correspond to the maximum amount of vertical and horizontal padding that can be added to the top and left of the rescaled view. We proceed analogously to sample a pair of final spatial coordinates $(C_{T-1}^{x},C_{T-1}^{y})$, given $(H_{T-1},W_{T-1})$.
[…]

### v2

@@ 4 Experiments
We train our model on over 15 million 24-frame clips from publicly-available online videos, in conjunction with standard training on Kubric. The resulting model is essentially a drop-in replacement for TAPIR (albeit with slightly larger computational requirements due to the extra layers). We evaluate on the TAP-Vid benchmark using the standard protocol.
@@ 4.1 Training datasets
We collected a video dataset from publicly accessible videos selected from categories that typically contain high-quality and realistic motion (such as lifestyle and one-shot videos). Conversely, we omitted videos from categories with low visual complexity or unrealistic motions, such as tutorial videos, lyrics videos, and animations. To maintain consistency, we exclusively obtained videos shot at 60fps. Additionally, we applied a quality metric by only considering videos with over 200 views. We removed the first and last 2 seconds of each video, as these often contain intros and outros with text or other overlays. From each video, we randomly sampled five clips, excluding those with overlay/watermarked frames, which were identified by checking the horizontal and vertical gradients and computing the pixel-wise median (similar to [9]). Furthermore, we expect the teacher signal will be more reliable on continuous shots due to temporal continuity; therefore, clips with shot boundary changes are detected and removed based on [5, 73, 61, 41] with additional accuracy improvements based on full-frame geometric alignment. In total, we generated 15 million clips for training.
@@ 4.2 Evaluation datasets
We rely on the TAP-Vid [12] and RoboTAP [62] benchmarks for quantitative evaluation; in all cases, we evaluate zero-shot on the entire benchmark, resizing to $256\times 256$ before evaluating according to the standard procedure [12]. This consists of five datasets: TAP-Vid-Kinetics which contains online videos of human actions and may include cuts [7]; TAP-Vid-DAVIS which is based on the DAVIS object tracking benchmark [49]; TAP-Vid-RGB-Stacking which contains synthetic tracks for videos of robotic manipulation which have little texture; and RoboTAP which contains real-world robotic manipulation videos [62], all of which include ground truth. Evaluation is performed by measuring occlusion accuracy (OA), $\mathbf{<\delta^{x}_{avg}}$ which measures the fraction of point estimates within a specified distance to the ground truth location, averaged across 5 thresholds, and Average Jaccard (AJ) which measures a combination of these two. There are two dataset querying “modes”: query first (q_first) uses the first visible point on each trajectory as a query, while strided uses every fifth point along the trajectory as a separate query. We also include qualitative evaluations on two robotics datasets without ground truth: RoboCAT-NIST, a subset of the data collected for RoboCat [6], and Libero [35], a dataset where point tracking has already proven useful for robotic manipulation [69]. See Appendix 0.D for details on these datasets and metrics.
@@ 4.3 Results
Table 2 shows performance under q_first mode. Here, we see that bootstrapping outperforms prior works by a wide margin on Kinetics; this is likely because TAPIR’s global search is more robust to large occlusions and cuts, which are more prominent in Kinetics. This search might harm performance in datasets like DAVIS with a stronger temporal continuity bias. Perhaps most impressive is the strong improvement in RoboTAP–over 5% absolute performance–despite RoboTAP looking very different from typical online videos. We see similar results for RGB-Stacking in Table 1. These two datasets have large textureless regions; such regions are challenging to track without object-aware priors, which are difficult to obtain from synthetic datasets.
Figure 3 shows qualitative examples of some cases where BootsTAPIR improves performance. We see improvements on examples where texture cues are ambiguous (e.g. the dark jacket and trousers) where prior knowledge of common object shape can improve performance, as well as points near object boundaries (e.g. the dog’s ears) where a model trained on synthetic data with different appearance may struggle to estimate the correct segmentation. We also note that BootsTAPIR improves on many cases where TAPIR marks a point as occluded even when it is still visible, such as the person’s arm. On RoboTAP, the model improves on occlusion estimation for the textureless gripper. It also deals well with large changes in scale as the gripper approaches the shoe, as well as shiny objects, both of which are less common in Kubric. Our project webpage https://bootstap.github.io/ includes video examples, which makes the improvements more obvious.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
@@
Figure 4: Comparison between TAPIR and BootsTAPIR on the real RoboCAT-NIST dataset. Due to the lack of ground truth, we show the TAPIR prediction and BootsTAPIR prediction in Rainbow tail style side-by-side. On NIST, BootsTAPIR works more consistently on location prediction. Particularly points that were originally predicted as occluded now can be visible.
@@
Figure 4 further illustrates improvements on the RoboCAT-NIST. Due to the lack of ground truth, sample a grid of points on the red pixels for RoboCAT-NIST. We display a few examples comparing the predicted tracks between the two models. As these are rigid objects, we expect the points to move consistently within each gear; deviations from this are errors. Due to the lack of texture on the gears and the nontrivial domain gap, the original TAPIR trained on Kubric works poorly here, with many jittery tracks and severe tracking failures. This is particularly bad for points that are close to occlusion or out of image boundary. The bootstrapped model fixes many of these failures: the tracks are much smoother and occlusion predictions become much more accurate. Results are comparable on Libero, although the motions there are more complicated and unsuitable for a static figure; see our project webpage for video visualizations.
4.4 Ablations
We focus on four main areas of ablation: data transformations, pseudolabel filtering approaches, training setup, and training data. To arrive at our final model, we performed ablations on a smaller-scale base setting with our best guesses at the optimal hyperparameter settings. This setting includes two components that we found could be removed without harming performance: It uses an additional mask on the occlusion loss, inspired by FixMatch [55], where any occlusion estimate that the teacher is uncertain about ($\max(\sigma(\hat{o}_{\mathcal{T}}[t]),1-\sigma(\hat{o}_{\mathcal{T}}[t]))<0.6$ is ignored in the loss. It uses a 3D-ConvNet backbone, which we find provides a slight improvement on DAVIS while harming performance on Kinetics (see Appendix 0.C), so we remove it for future compatibility with causal TAPIR models. Finally, base also halves the batch sizes (and proportionally halves the learning rate), and also halves the number of training steps. We report Average Jaccard on DAVIS using the strided mode and on Kinetics using the q_first mode.
Data transformations.
We first investigate the effect of the transformations we apply on inputs and outputs in this setting. We respectively ablate: the use of random JPEG augmentations to enforce invariance to various factors of variation (denoted by base-no-augm); the use of framewise affine transformations on inputs and outputs to enforce equivariance with spatial transformations (denoted by base-no-affine). We also investigate sampling the student queries: recall that in our typical setup, we sample the student query from a distribution which places probability 0.5 on the original teacher query point, and 0.5 on a uniform distribution across visible points. In base-same-queries, we always use the teacher’s query for the student, and in base-uniform, we sample from a purely uniform distribution. We report the results for each ablation in Table 3(a) (a). We observe that removing JPEG somewhat harms metrics, especially on Kinetics. In contrast, when ablating affine transformations, we find that performance drops massively across metrics, suggesting overfitting. Finally, we find that using different query points improves performance compared with base-same-queries, leading to more accurate position predictions in particular ($\mathbf{<\delta^{x}_{avg}}$ increases from 77.5 to 77.9 on DAVIS strided and from 66.8 to 67.7 on Kinetics q_first). Note, however, on DAVIS in particular, this improvement depends on sampling the original teacher query point more often than the others.
Table 3: Ablations of model hyperparameters, including (a) ablations of the data transformations and query point strategies, (b) comparisons of the pseudolabel filtering approaches, (c) ablations of the training setup, including the stop gradient, and (d) ablations of the dataset. We report Average Jaccard (AJ) across all experiments.
DAVIS Kinetics
Method strided q_first
base 65.8 54.4
base-no-augm 65.7 53.5
base-no-affine 54.4 44.7
base-same-queries 65.6 53.2
base-uniform 65.6 54.3
(a) Data transformations.
DAVIS Kinetics
Method strided q_first
base 65.8 54.4
base-no-filtering 65.9 54.1
base+cycle 66.1 54.3
(b) Pseudolabels filtering. base filters occlusion loss terms based on teacher confidence.
DAVIS Kinetics
Method strided q_first
Full model 66.2 54.6
full-kubric-only 65.0 52.7
base 65.8 54.4
siamese 49.8 29.6
(c) Training setup.
DAVIS Kinetics
Method strided q_first
base 65.8 54.4
2-frame clips 64.3 50.5
6-frame clips 63.7 50.9
1% of real data 66.2 54.0
(d) Training Data.
Pseudolabel filtering.
We next consider the effectiveness of filtering possibly-incorrect teacher tracks and points, with results in Table 3(b) (b). base-no-filtering removes the filtering that base uses on the occlusion confidence score, which makes little difference in performance on DAVIS, but degrades performance on Kinetics. base+cycle uses cycle-consistency criterion from our full model instead and performs slightly better on DAVIS. These results suggest that correctly removing bad teacher tracks remains an open problem.
Training setup.
Table 3(c) (c) shows ablations of the overall training setup. In particular, training for longer with a higher capacity model can improve results, so full-kubric-only uses an identical training setup to our full model, but removes self-supervised training, instead simply training on Kubric for longer. We see competitive performance, although self-supervised training still improves by 1.2% on DAVIS and almost 2% on Kinetics. siamese shows the effect of removing the EMA and stop-gradient and instead backpropping to both student and teacher models (in this case, using the base setting): we note that performance on real-world datasets collapses as the model finds trivial shortcuts.
Training data.
We ablate two questions regarding the dataset. Prior work has argued that simple semi-supervised learning for optical flow performs poorly [31, 39]; we hypothesize that more temporal context may be the key ingredient to change this story. To validate this, we re-ran our algorithm using 2- and 6-frame clips from our full dataset. In Table 3(d) (d), we see that this indeed performs poorly, possibly because the extra frames allow the teacher model to correct more errors. Interestingly, we also tried training on a 1% subset of the data, and found that this harms performance on Kinetics but actually improves it on DAVIS. It’s possible that the algorithm begins overfitting to the data, but this may be useful for clean data like DAVIS. Regardless, it suggests this algorithm can be effective even in situations where less data is available.
5 Higher Resolution and Public Release
The publicly released version of BootsTAPIR contains a fix for a minor bug that was present in our previous versions of TAPIR. Specifically, recall that the data augmentations used for the Kubric dataset include a random axis-aligned crop. The image cropping mechanism was not pixel-aligned with the transforms used for points, leading to an almost imperceptible error in the track locations. Fixing this bug leads to an improvement in performance for the original TAPIR model, but surprisingly has relatively little effect on BootsTAP performance. However, we find that the reason isn’t because BootsTAP compensates for the bug, but rather, because the bug creates a bias toward tracking foreground objects (the tracks tend to be slightly expanded relative to the underlying objects). We find we can replicate this bias by altering query points that are very near occlusion edges (1 pixel away) to track the foreground object rather than the background, which we call the “snap to occluder” technique. See Appendix 0.B.4 for details.
To further tune performance, we also trained on higher-resolution clips, and also longer clips, as we find these improve generalization for real-world applications with longer or higher-resolution videos. To implement this, we add more ‘tasks’ with different data shapes, using the same multi-task framework (i.e., separate optimizers) as described above. Specifically, one extra task uses $512\times 512$ Kubric clips (24 frames), trained using the same losses. We use the hierarchical refinement approach described in the original TAPIR paper, wherein the initialization and one refinement pass is performed at $256\times 256$, and then a further refinement pass is performed at $512\times 512$. We also use an analogous high-resolution self-supervised task, which also uses 24-frame, $512\times 512$ videos from the same real-world dataset. Finally, we add 150-frame, $256\times 256$ videos, this time at 30 frames per second.
Table 4: Comparison of performance on the TAP-Vid datasets for the released version of BootsTAPIR. Fix refers to the bugfix to coordinates. Snap refers to the snap-to-occluder bias in the training data. Data refers to extra training data which has longer clips and higher resolution.
Kinetics DAVIS RGB-Stacking
Method AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA
CoTracker [29] 57.3 70.6 87.5 64.8 79.1 88.7 65.9 80.6 85.0
Tuned TAPIR+fix 60.4 72.5 88.6 64.7 76.6 89.6 70.0 82.1 89.9
Tuned TAPIR+fix+snap 59.6 71.9 88.8 63.8 75.9 89.9 67.1 80.3 87.8
BootsTAPIR 61.4 74.2 89.7 66.2 78.1 91.0 72.4 83.1 91.2
BootsTAPIR+fix 61.4 74.5 89.4 66.5 78.5 90.8 75.9 85.7 93.8
BootsTAPIR+fix+snap 61.3 74.7 89.1 67.1 78.9 91.2 75.7 85.7 93.4
BootsTAPIR+fix+snap+data 62.5 74.8 89.5 67.4 79.0 91.3 77.4 86.7 93.2
Kinetics DAVIS RoboTAP
Method AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA
CoTracker [29] 48.7 64.3 86.5 60.6 75.4 89.3 54.0 65.5 78.8
Tuned TAPIR+fix 53.3 66.0 85.1 58.9 71.6 86.4 67.3 78.4 90.0
Tuned TAPIR+fix+snap 52.5 65.3 85.5 58.3 71.1 87.7 66.4 77.3 90.5
BootsTAPIR 54.6 68.4 86.5 61.4 73.6 88.7 64.9 80.1 86.3
BootsTAPIR+fix 54.7 68.5 86.3 61.6 74.1 89.0 65.7 80.5 87.2
BootsTAPIR+fix+snap 54.5 68.8 86.3 61.8 74.3 89.1 63.5 81.1 84.2
BootsTAPIR+fix+snap+data 55.8 68.8 86.6 62.4 74.6 89.6 69.2 81.3 89.5
Tables 4 and 5 show our results. Note that CoTracker implemented its own data augmentation algorithms and is not affected by the same bug. We see that “snap to occluder” harms TAPIR performance, but improves BootsTAP performance. One possible interpretation is that the snapping is compensating for a particular bias in the bootstrapping toward tracking background. This may be because background is easier to track, especially relative to thin objects. In a bootstrapping framework, the model’s reliable predictions that follow background become self-reinforcing, whereas unreliable predictions for thin foreground objects, are not. Therefore, they tend to get lost over time. Finding more principled solutions to this issue is an interesting area for future work.
The extra training data, however, leads to a non-trivial boost in performance. Note that, for tables 4 and 5, all evaluation videos are still at $256\times 256$, and unlike many prior methods we do not upsample them before creating the feature representation. To assess the impact of increased evaluation resolution, we also performed evaluation on $512\times 512$ videos. Table 6 shows results. We see that performance improves by 1.2% on Kinetics and 2.8% on DAVIS, the best reported performance on this dataset by a wide margin. Surprisingly, we found that further resolution at test-time did not improve results, suggesting another interesting area for future work.
Table 6: Comparison of performance for high-resolution setting.
Kinetics DAVIS
Method AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA
BootsTAPIR 62.5 74.8 89.5 67.4 79.0 91.3
BootsTAPIR+hires 63.7 76.0 88.4 70.2 81.2 91.2
RoboTAP [62] pointed out that point tracking can be very useful in an online setting, e.g. when used as a signal to control agents in real time. It remains straightforward to extend BootsTAPIR to the online setting: the only temporal dependency of the model is in the 1D convolutions in the iterative refinements, so these can be directly converted into causal convolutions to create a causal model. We trained this model using the full training setup for the release model, including the extra high-resolution, long-clip data. Table 7 shows results. We see an overall 4.6% improvement on Kinetics and a 3.0\$ improvement on DAVIS, in both cases using the query-first evaluation procedure.
Table 7: Causal model performance.
Kinetics q_first DAVIS q_first
Method AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA
Causal TAPIR 51.5 64.4 85.2 56.7 70.2 85.7
Causal BootsTAPIR 55.1 67.5 86.3 59.7 72.3 86.9
We also perform experiments on the point tracks in the Perception Test [48] validation set, a challenging dataset of point tracks annotated on videos of unusual situations filmed by participants. Results are shown in Table 8; we see a similar magnitude gap over prior results.
Table 8: Performance on Perception Test relative to TAPIR.
Method Overall Static Dynamic
TAPIR 55.7 57.4 46.3
BootsTAPIR 59.6 61.3 49.7
As a final note, we performed informal benchmarking of our model using an A100 and the latest JAX compiler. We found that after compilation, BootsTAPIR can perform inference of $10,000$ points on a $256\times 256$, 50-frame video in 5.6 seconds. Furthermore, the causal model can track 400 points on a $256\times 256$ video at 30.1 frames per second.
@@ Appendix 0.A Summary of the approach
We summarize notation and computation of our self-supervised loss in Algorithm 1.
Algorithm 1 BootsTAP self-supervised loss. Notation:
@@
$\mathcal{V}\mapsto\mathcal{D}_{\mathcal{V}}$ – mapping that maps a set of points $\mathcal{V}$ to a distribution $\mathcal{D}_{\mathcal{V}}$ over $\mathcal{V}$
@@
    sample $Q_{2}=(q_{2},t_{2})\sim\mathcal{D}_{\mathcal{V}}$.
@@ Appendix 0.B Implementation details
0.B.1 Distribution over transformations
We design transformations of the inputs to enforce equivariance of the predictions with realistic spatial transformations. At a high level, we intend for the transformations to mimic the effects of additional, simple and plausible camera motion and zooming on the video. Hence, our transformations should vary smoothly in time; they should cover a reasonable ratio of the original video content; and aspect ratio should be roughly preserved.
We define a family of frame-wise affine transformations that has these properties, and a procedure to sample these randomly. Essentially, we sample top-left crop coordinates and crop dimensions for each frame in the video; where coordinates and dimensions are computed as interpolations between values sampled for the start and end frames from a distribution that achieves the desired coverage and aspect ratios.
More formally, we first sample a pair of spatial dimensions $(H_{0},W_{0})$ for the start frame as follows. We sample an area $A$ uniformly over $[0.6,1.0]$. Next we sample values $a^{1},a^{2}\sim\mathcal{U}([A,1])$ and derive random height value by averaging them $h=\frac{a^{1}+a^{2}}{2}$ and width value $w=\frac{A}{h}$; and finally, we multiply these values by the input’s original shape $(H,W)$. This gives us a pair of spatial dimensions biased towards aspect ratios close to 1, and covering an area between $60\%$ and $100\%$ of the original input. We proceed the same way to sample a pair of spatial dimensions$(H_{T-1},W_{T-1})$ for the end frame.
Next, we uniformly sample a pair of top-left corner coordinates $(C_{0}^{x},C_{0}^{y})$ for the start frame, such that a crop of dimensions $(H_{0},W_{0})$ can be extracted within the frame. We proceed the same way to sample a pair of spatial coordinates $(C_{T-1}^{x},C_{T-1}^{y})$, given $(H_{T-1},W_{T-1})$, for the end frame.
We then interpolate linearly on one hand between the start and end spatial dimensions; and on the other hand between the start and end top-left corner coordinates. Let $t\in\{0,...,T-1\}$ be a frame index. Calling $\alpha_{t}=\frac{t}{T-1}$, we define: